The a umulated amount that you will re eive after 52 months if you deposit R7 300 into an a ount whi h earns 9,7% interest per year, ompounded every se ond month, is

[1] R7 825,36.

[2] R10 368,43.

[3] R11 076,73.

[4] R8 388,53.

[5] none of the above.


1
Expert's answer
2021-06-27T04:05:23-0400

Given that:

"FV=P(1+\frac{r}{m})^{nm}"

Where FV is the future value,

PV= R7300

r=9.7%

n="4\frac{1}{3}" years

m=6


"\therefore FV=7300(1+\frac{0.097}{6})^{6×4\frac{1}{3}}"

"FV=R11076.73"


Option [3] is the right answer